OpenStudy (anonymous):

A quadratic equation is shown below: 3x2 - 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the discriminant. Show your work. (3 points) Part B: Solve 9x2 + 3x - 2 = 0 by using an appropriate method. Show the steps of your work and explain why you chose the method used. (4 points) Part C: Solve 3x2 - 10x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (solomonzelman):

Do you know what a discriminant is ?

OpenStudy (anonymous):

no I dont

OpenStudy (solomonzelman):

Have you ever used a quadratic formula ? If so, then you should be familiar with this, \(\LARGE\color{black}{ x= \frac{-b±\sqrt{\color{blue}{b^2-4(a)(c)}}}{2a} }\) . The discriminant is the blue part.

OpenStudy (anonymous):

ohhhh yes I remember

OpenStudy (solomonzelman):

you will know that if b²-4ac is equal to a negative number (when you plug your values in) then you have no (real) solutions. IF your discriminant - the b²-4ac is a zero, then you will have 1 solution. And if it is equal to some positive number (when you plug your a,b,c into it) then you get 2 soltuions. (and if the discriminant, the b²+4ac is a perfect square, (such as 1, 4, 9, 16, 25, 36 and on ) then you can factor it )

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \color{blue}{3}x^2 \color{green}{-16}x + \color{red}{3} = 0}\) Compare to \(\normalsize\color{black}{ \color{blue}{a}x^2 +\color{green}{b}x + \color{red}{c} = 0}\)

OpenStudy (anonymous):

okay I will

OpenStudy (solomonzelman):

Can you find the discriminant ? Tell me what you get

OpenStudy (anonymous):

I don't tjhink im doing it right

OpenStudy (solomonzelman):

you say that you don't think you are dong this right, but can I please see WHAT you are doing, to verify ?

OpenStudy (solomonzelman):

Can you find what b² - 4ac is going to be equal to in YOUR case, or need help with that ?

OpenStudy (anonymous):

(x+6)(x-3)?

OpenStudy (solomonzelman):

You need to just find the discriminant. Don't solve it yet:)

OpenStudy (anonymous):

\[10\pm \sqrt{(-10)^{2}}-4(3)(2)\]

OpenStudy (anonymous):

2(3) on the bottom

OpenStudy (solomonzelman):

Which equation are you doing? I was referring to the PART A, where you have 3x²-16x+2=0 And the discriminant is `(-16)² - 4(3)(2)` `256 - 24` `232`

OpenStudy (anonymous):

ohh im sorry

OpenStudy (anonymous):

so that was for part a?

OpenStudy (solomonzelman):

you get that the discriminant is 232, and when the discriminant is positive you will have 2 solutions.

OpenStudy (anonymous):

yupp

OpenStudy (solomonzelman):

I got disconnected. Durn it !

OpenStudy (anonymous):

its okay(:

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ 3x^2 - 10x + 2 = 0}\) use the quadratic formula, because you can't factor since the discriminant is not a perfect square (but it is a 74), and completing the square isn't good, because you already starting with an odd coefficient in front of x².

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ x=\frac{-(-10)±\sqrt{(-10)^2-4(3)(2)}}{2(3)} }\) go for it.